OpenStudy (anonymous):

ln ( x^5 sqroot x-1 / 5x+3 ) expand

myininaya (myininaya):

\[\ln(\frac{x^5 (x-1)^\frac{1}{2}}{5x+3})\]?

OpenStudy (anonymous):

yes

myininaya (myininaya):

do you recall \[\ln(\frac{a}{b})=\ln(a)-\ln(b)\]

myininaya (myininaya):

\[\ln(x^5(x-1)^\frac{1}{2})-\ln(5x+3)\]

myininaya (myininaya):

do you also recall \[\ln(ab)=\ln(a)+\ln(b)\]

myininaya (myininaya):

\[\ln(x^5)+\ln((x-1)^\frac{1}{2})-\ln(5x+3)\]

myininaya (myininaya):

do you recall \[\ln(y^r)=r \ln(y)\]

myininaya (myininaya):

\[5 \ln(x)+\frac{1}{2}\ln(x-1)-\ln(5x+3)\]

OpenStudy (anonymous):

oh yes the properties, sorry i forgot that from precalc, and i was trying help someone